OpenStudy (anonymous):

Let b = 10. What is the value of 6b – 21?

OpenStudy (anonymous):

help plz

OpenStudy (kendricklamar2014):

\[b = 10 ~ ~ 6*10-21\]

OpenStudy (kendricklamar2014):

Now you do:\[6*10 = 60 - 21\]

OpenStudy (kendricklamar2014):

What is: \[60-21 = ?\] @deadpoolheex

OpenStudy (anonymous):

60 - 21 = 39

OpenStudy (kendricklamar2014):

Yes, 60 - 21 = 39

OpenStudy (anonymous):

so the answer is 39 right

OpenStudy (kendricklamar2014):

Right

OpenStudy (anonymous):

thanks

OpenStudy (kendricklamar2014):

No Problem